OpenStudy (anonymous):

15. A $500 investment earns interest at 6% per year compounded quarterly for 8 years. How much interest will be earned in the final year of the 8 year term?

OpenStudy (anonymous):

So in seven years there are 7 * 4 = 28, if the annual interest is 6%, the quarterly interest should then be 1.5%. So in each quarter your money is multiplied by 1.015 so after the first quarter you will have 500 * 1.015 = 507.5 dollars. So multiplying the sum 28 times with 1.015 should net you 500 * 1.015^28 = 758.6110900487 dollars. If we add 4 more quarters we will have 32 quarters which net 805.1621600775935 dollars, subtracting the two numbers will give us the answer: 46.55107002889349 dollars are earned in the final year of the 8 year term. If I understand compound interest correctly of course.

OpenStudy (anonymous):

why did you do 7 years tho i dont understand that part i mean cant u just find the value after the eight year and then find interest using I = A - P